# Precise Measurements of the LMC Bar’s Geometry With Gaia DR3 and a Novel Solution to Crowding Induced Incompleteness in Star Counting

### v1

@@ 5 The Geometry of the LMC’s Bar
In this section, we apply the techniques discussed in the previous section to measure the LMC bar’s properties from our complete red clump star catalog. We shall analyze the bar properties in the LMC cartesian coordinate system, to obtain parameters in physical units.
@@ 5.1 Bar Position Angle
We illustrate this position angle in Figure 6. The position angle obtained from Fourier decomposition aligns very well with the visual expectation, affirming the reliability of our method. Our measured bar position angle implies that the LMC’s bar is misaligned with the line of nodes of the disk ($\sim 150^{\circ}$) (Choi et al., 2018a) by around 30.
@@
Figure 6: Various estimates for the center of mass of the LMC’s disk. We depict the dynamical center of mass for the completeness corrected and incomplete samples using the blue star and blue triangle respectively. The magenta “K” depicts the LMC kinematic center found by C22, the brown “P” the LMC photometric center found by van der Marel (2001), and the red “H” the HI kinematic center found by Kim et al. (1998). We find that incompleteness significantly affects the determination of the dynamical center of mass. The dotted black line depicts the position angle of the LMC’s bar calculated using Fourier decomposition. The remarkable agreement between the inferred phase angle and the visual expectation based on the location of the bar (shown by the black ellipse) confirms the reliability of the Fourier decomposition technique. In these coordinates, north is towards the top and east is towards the left.
@@ 5.2 Bar Strength and Radius
It is well known that the LMC’s bar is lopsided (van der Marel, 2001; Jacyszyn-Dobrzeniecka et al., 2016), which is likely due to a spiral arm attached to the western end of the bar. Asymmetries such as lopsidedness appear as odd m terms of the Fourier series (Pardy et al., 2016; Łokas, 2021), which are orthogonal to the m = 2 term. Indeed the m = 2 term has been successfully used to measure bar properties in interacting/distorted barred galaxy simulations (e.g. Gerin et al., 1990; Berentzen et al., 2003). Hence, we do not expect the lopsided nature of the LMC’s bar to affect the inference of the bar strength and radius from the m = 2 amplitude and phase.
@@ 5.3 Bar Axis-Ratio and Offset
Figure 8: Illustration of the offset bar of the LMC. The left panel shows the Gaia DR3 red clump disk, corrected for completeness. The right panel shows the LMC disk in the B12 simulation, where the SMC has undergone a recent direct collision with the LMC. The solid ellipse is the iso-contour corresponding to the bar, and the dashed ellipse is the iso-contour corresponding to the outer disk. The star denotes the center of the iso-ellipse of the bar, and the filled circle denotes the center of the iso-ellipse of the outer disk. The two centers do not coincide, which indicates an offset bar. The offset is seen in the simulated disk as well, however, it is larger relative to the observations.
@@
We compile all the measurements of this section, for both the completeness corrected and incomplete samples in Table 3. We compare our measurements of the bar parameters with a subset of literature values reported in Table 1. In section 6.3, we shall place our measurements in context with the population of barred galaxies in the local universe.
As mentioned in Section 1, the LMC’s bar is likely in a different plane relative to the overall LMC disk (Subramanian & Subramaniam, 2013; Choi et al., 2018a). A different inclination of the bar region with respect to the outer disk can cause a projection effect in our estimate of the bar length and the bar offset. However, we expect this effect to be small, since the average tilt of the LMC’s bar plane relative to the disk plane is around 10 degrees, which corresponds to a projection bias factor of cos(10) = 0.98, which is very close to unity. Thus, the expected projection bias is within our errors for the bar measurements. In addition, the LMC’s inner regions might have a different position angle relative to the LMC’s outer regions (Choi et al., 2018a; Dhanush et al., 2024). However, the Fourier decomposition technique adopted in this work averages over the azimuthal coordinate. Thus the global position angle at which we choose to orient our coordinate system will not significantly affect our results as long as we ensure that the position angle is constant in the bar region.
@@ 6.1 Comparison of our Measured Bar Parameters with Literature
We find that the LMC bar radius measurement for the completeness corrected sample ($2.13^{+0.03}_{-0.04}$ kpc) is close to the measurement of JA24 (2.3 kpc). JA24 used the dynamical model by Dehnen et al. (2023) based on the stellar continuity equation, which yields both the bar pattern speed as well as the bar radius. This dynamical model identifies the bar as a collection of stars with a roughly constant bi-symmetric phase, which is similar to our approach in principle. We could not find error estimates in JA24, thus it is difficult to compare their measurement with ours in a statistical sense. However, given the similarities between our methods, it is reasonable that the two measurements are within 0.5 kpc.
@@
We find the LMC’s bar position angle measurement (121.26 $\pm$ 0.21) to be consistent with previous estimates like Zhao & Evans (2000); van der Marel (2001) (see Table 1). However, our position angle is significantly different from C18b (154.18). We attribute this difference to another subtlety of the combined bar $+$ disk model. They mention that when the radius of the bar reaches the maximum value allowed by the prior, the model tries to fit the surrounding stars by becoming boxier. Because of this, we anticipate that the model is including a part of the disk stars in the bar, which biases the inferred position angle of the bar.
We find the LMC’s bar axis ratio measurement (0.54 $\pm$ 0.03) to be larger than the estimates of de Vaucouleurs & Freeman (1972); Zhao & Evans (2000); Choi et al. (2018b), and on the higher side of the range described by van der Marel (2001) (0.33 - 0.57). This indicates that the LMC bar might be thicker than previously thought.
We find the LMC’s bar offset (0.76 $\pm$ 0.01 kpc) to be larger compared to the estimate of van der Marel (2001) (0.4 kpc), but closer to the estimate of Zhao & Evans (2000) (0.87 kpc). Accurately measuring the bar offset is important for the dynamical modelling of the LMC’s bar in presence of the SMC’s interactions (see next section).
@@ 6.2 Comparison With a Hydrodynamic Model of a Recent LMC-SMC Collision
To obtain insights on the evolutionary history of the LMC through its measured bar properties, we compare our measurements with the hydrodynamic model of Besla et al. (2012) (hereafter B12). B12 modeled the interaction history of the LMC and SMC over the past 6-7 Gyr, including a MW infall for the past 1 Gyr. B12 presented two models - Model 1 and Model 2. In Model 1, the SMC and LMC remain far from each other, with the closest passage being $\geq 20$ kpc. In Model 2, the SMC undergoes a direct collision with the LMC’s disk (impact parameter $\sim 2$ kpc) around $100$ Myr ago. Model 2 reproduces the observed structure and kinematics of the LMC significantly better compared to Model 1 (Besla et al., 2016; Choi et al., 2018a; Zivick et al., 2019; Choi et al., 2022), thus favoring a direct collision scenario. Moreover, the bar in Model 1 is not consistent with observations (B12). It does not show any offset with respect to the disk and does have a prominent signature in gas. Hence, we compare our measurements to B12 Model 2. Below, we provide a brief summary of the B12 simulation setup.
@@
Note that the simulations of the LMC-SMC-MW interaction history used to compare against our observations (B12, Model-2) do not reproduce some key observables of the LMC-SMC system, like the SMC’s present day Galactocentric position and velocity and the LMC’s disk inclination. However, our goal is primarily to show that the observed LMC bar structure is consistent with a scenario where the SMC has undergone a recent collision with the LMC. The simulations are well-suited to study this scenario in general. New simulations that are better matched to observations are needed for the purpose of an in-depth comparative study, which is work in progress.
@@ 6.3 Placing the LMC in Context with Other Barred Galaxies
7 Other Applications of the Completeness Correction Framework
@@ 8 Conclusions
For the first time, we find the LMC bar’s strength, as quantified by the Fourier bi-symmetric amplitude, to be $S_{bar}=0.27$. This values is significantly above the threshold ($S_{bar}=0.20$) used to define barred galaxies, meaning the LMC is a traditional barred galaxy.
Through this work, we have presented a reliable geometric framework that can accurately determine multiple bar parameters, without relying on assumptions of the mass profile of the bar and disk. We find a general agreement between the bar parameters measured from our framework and literature measurements that are based on similar methods (e.g. Jiménez-Arranz et al., 2024).
@@
Himansh Rathore would like to acknowledge interesting discussions with Mathieu Renzo, Peter Behroozi, Leandro Beraldo-e-Silva, David Sand and Dennis Zaritsky regarding the LMC’s bar as well as the completeness calibration framework. We would like to thank the anonymous referee for providing insightful comments that improved the quality as well as clarity of the paper. This work utilized the Puma and ElGato High Performance Computing clusters at the University of Arizona. Himansh Rathore and Gurtina Besla are supported by NSF CAREER AST 1941096. This research uses services or data provided by the Astro Data Lab, which is part of the Community Science and Data Center (CSDC) Program of NSF NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the U.S. National Science Foundation. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. We respectfully acknowledge the University of Arizona is on the land and territories of Indigenous peoples. Today, Arizona is home to 22 federally recognized tribes, with Tucson being home to the O’odham and the Yaqui. Committed to diversity and inclusion, the University strives to build sustainable relationships with sovereign Native Nations and Indigenous communities through education offerings, partnerships, and community service.
@@ Appendix A Gaia Data Selection
where $\omega$ is the parallax in mas; $\mu_{\alpha\ast}$ and $\mu_{\delta}$ are the proper motions in Right Ascension (RA; $\alpha$, $\alpha^{\ast}=\alpha\cos\delta$) and declination (DEC; $\delta$) in mas yr-1 respectively; $\rho$ is the galactic radius from the LMC kinematic center $(\alpha_{0},\delta_{0})=(80.443^{\circ},-69.272^{\circ})$ as determined in C22. We use the following query in the Gaia query interface:
@@ Appendix B Coordinate Systems
The LMC cartesian coordinates can be obtained by the following transformation:

### v2

@@ 5 The Geometry of the LMC’s Bar
In this section, we apply the techniques discussed in the previous section to measure the LMC bar’s properties from our complete red clump star catalog. We shall analyze the bar properties in the LMC in-plane cartesian coordinate system, to obtain parameters in physical units.
@@ 5.1 Bar Position Angle
We illustrate this position angle in Figure 6. The position angle obtained from Fourier decomposition aligns very well with the visual expectation, affirming the reliability of our method. Our measured bar position angle implies that the LMC’s bar is misaligned with the line of nodes of the disk ($\sim 150^{\circ}$) (Choi et al., 2018a) by around 30, which is consistent with JA24.
@@
Figure 6: Various estimates for the center of mass of the LMC’s disk. We depict the stellar dynamical center of mass for the completeness corrected and incomplete samples using the blue star and blue triangle respectively. The magenta “K” depicts the LMC stellar kinematic center found by C22, the brown “P” the LMC photometric center found by van der Marel (2001), and the red “H” the HI kinematic center found by Kim et al. (1998). We find that incompleteness significantly affects the determination of the stellar dynamical center of mass. The dotted black line depicts the position angle of the LMC’s bar calculated using Fourier decomposition. The remarkable agreement between the inferred phase angle and the visual expectation based on the location of the bar (shown by the black ellipse) confirms the reliability of the Fourier decomposition technique. In these coordinates, north is towards the top and east is towards the left.
@@ 5.2 Bar Strength and Radius
We do not expect the bar offset to significantly affect our analysis, since we have ensured accurate centering on the bar, as described in section 4.2. In addition to the offset, it is well known that the LMC’s bar is also lopsided (more mass being present on one end of the bar relative to the other). The lopsidedness is likely due to a spiral arm attached to the western end of the bar (van der Marel, 2001; Jacyszyn-Dobrzeniecka et al., 2016). Bar Lopsidedness manifests as odd-m terms of the Fourier series (Pardy et al., 2016; Łokas, 2021). In the completeness corrected sample, we find a significant $A_{1}$ component in the bar region after centering on the bar, confirming its lopsided nature. Since the odd-m terms of the Fourier series (equations 11 and 12) are orthogonal to the even-m terms, we do not expect the bar lopsidedness to significantly affect our measurements of bar strength and radius from the m = 2 term. Indeed, the m = 2 term has been successfully used to measure bar properties in interacting/distorted barred galaxy simulations which show asymmetries such as offset and lopsidedness, after accurate centering (e.g. Gerin et al., 1990; Berentzen et al., 2003).
@@ 5.3 Bar Axis-Ratio and Offset
Figure 8: Illustration of the offset bar of the LMC. The left panel shows the Gaia DR3 red clump disk in the LMC in-plane cartesian coordinates, corrected for completeness. The right panel shows the LMC disk in the B12 simulation, where the SMC has undergone a recent direct collision with the LMC. The solid ellipse is the iso-contour corresponding to the bar, and the dashed ellipse is the iso-contour corresponding to the outer disk. The star denotes the center of the iso-ellipse of the bar, and the filled circle denotes the center of the iso-ellipse of the outer disk. The two centers do not coincide, which indicates an offset bar. The offset is seen in the simulated disk as well, however, it is larger relative to the observations.
@@
We compile all the measurements of this section, for both the completeness corrected and incomplete samples in Table 3. We compare our measurements of the bar parameters with a subset of literature values reported in Table 1 in section 6.1 and with a numerical simulation of the LMC-SMC-MW interaction history in section 6.2. In section 6.3, we shall place our measurements in context with the population of barred galaxies in the local universe.
As mentioned in Section 1, the LMC’s bar is likely in a different plane relative to the overall LMC disk (Subramanian & Subramaniam, 2013; Choi et al., 2018a). A different inclination of the bar region with respect to the outer disk can cause a projection effect in our estimate of the bar length and the bar offset. However, we expect this effect to be small, since the average tilt of the LMC’s bar plane relative to the disk plane is around 10, which corresponds to a projection bias factor of $\cos(10^{\circ})=0.98$, which is very close to unity. Thus, the expected projection bias is within our errors for the bar measurements. In addition, the LMC’s inner regions might have a different position angle relative to the LMC’s outer regions (Choi et al., 2018a; Dhanush et al., 2024). However, the Fourier decomposition technique adopted in this work averages over the azimuthal coordinate. Thus the global position angle at which we choose to orient our coordinate system will not significantly affect our results as long as we ensure that the position angle is constant in the bar region.
@@ 6.1 Comparison of our Measured Bar Parameters with Literature
We find that the LMC bar radius measurement for the completeness corrected sample ($2.13^{+0.03}_{-0.04}$ kpc) is close to the measurement of JA24 (2.3 kpc). JA24 used the dynamical model by Dehnen et al. (2023) based on the stellar continuity equation, which yields both the bar pattern speed as well as the bar radius. This dynamical model identifies the bar as a collection of stars with a roughly constant bi-symmetric phase, which is similar to our approach in principle. We could not find error estimates in JA24, thus it is difficult to compare their measurement with ours in a statistical sense. However, given the similarities between our methods, it is reasonable that the two measurements are within 0.5 kpc. Also, JA24 estimated the amplitude of the m = 2 component of the velocity field in the LMC’s central region. The peak radius of the velocity m = 2 amplitude (Figure 7 of JA24) is consistent with the peak radius of the stellar density m = 2 amplitude (Figure 7) within 0.5 kpc. Our bar radius measurement is also consistent with Kacharov et al. (2024), who infer a radius of 2.2 kpc with Schwarzschild Orbit Modeling (Schwarzschild, 1979) of the LMC’s inner regions.
@@
We find the LMC bar’s position angle measurement (121.26 $\pm$ 0.21) to be consistent with previous estimates like Zhao & Evans (2000); van der Marel (2001) (see Table 1). However, our position angle is significantly different from C18b (154.18). We attribute this difference to another subtlety of the combined bar $+$ disk model. They mention that when the radius of the bar reaches the maximum value allowed by the prior, the model tries to fit the surrounding stars by becoming boxier. Because of this, we anticipate that the model is including a part of the disk stars in the bar, which biases the inferred position angle of the bar.
We find the LMC bar’s axis ratio measurement (0.54 $\pm$ 0.03) to be larger than the estimates of de Vaucouleurs & Freeman (1972); Zhao & Evans (2000); Choi et al. (2018b), and on the higher side of the range described by van der Marel (2001) (0.33 - 0.57). This indicates that the LMC bar might be thicker than previously thought.
We find the LMC bar’s offset (0.76 $\pm$ 0.01 kpc) to be larger compared to the estimate of van der Marel (2001) (0.4 kpc), but closer to the estimate of Zhao & Evans (2000) (0.87 kpc). Accurately measuring the bar offset is important for the dynamical modelling of the LMC’s bar in presence of the SMC’s interactions (see next section).
@@ 6.2 Comparison With a Hydrodynamic Model of a Recent LMC-SMC Collision
To obtain insights on the evolutionary history of the LMC through its measured bar properties, we compare our measurements with the hydrodynamic model of Besla et al. (2012) (hereafter B12). B12 modeled the interaction history of the LMC and SMC over the past 6-7 Gyr, including a MW infall for the past 1 Gyr. B12 presented two models - Model 1 and Model 2. In Model 1, the SMC and LMC remain far from each other, with the closest passage being $\geq 20$ kpc. In Model 2, the SMC undergoes a direct collision with the LMC’s disk (impact parameter $\sim 2$ kpc) around $100$ Myr ago. Model 2 reproduces the observed structure and kinematics of the LMC significantly better compared to Model 1 (Besla et al., 2016; Choi et al., 2018a; Zivick et al., 2019; Choi et al., 2022), thus favoring a direct collision scenario. Moreover, the bar in Model 1 is not consistent with observations. It does not show any offset with respect to the disk and does have a prominent signature in gas. Hence, we compare our measurements to B12 Model 2. Below, we provide a brief summary of the B12 simulation setup.
@@
Note that the simulations of the LMC-SMC-MW interaction history used to compare against our observations (B12, Model-2) do not reproduce some key observables of the LMC-SMC system, like the SMC’s present day Galactocentric position and velocity and the LMC’s disk inclination. However, our goal is primarily to show that the observed LMC bar structure is consistent with a scenario where the SMC has undergone a recent collision with the LMC. The simulations are well-suited to study this scenario in general. New simulations that are better matched to observations are needed for the purpose of an in-depth comparative study, which is work in progress. Future work also involves comparison of the observed bar properties with a library of simulated LMC-like disks under a variety of interaction scenarios for LMC and SMC like galaxies (e.g. Jiménez-Arranz et al., 2024a, KRATOS simulations).
@@ 6.3 Placing the LMC in Context with Other Barred Galaxies
6.4 Other Applications of the Completeness Correction Framework
@@ 7 Conclusions
For the first time, we find the LMC bar’s strength using the stellar density field, as quantified by the Fourier bi-symmetric amplitude, to be $S_{bar}=0.27$. This value is significantly above the threshold ($S_{bar}=0.20$) used to define barred galaxies, meaning the LMC is a traditional barred galaxy.
Through this work, we have presented a reliable geometric framework that can accurately determine multiple bar parameters, without relying on assumptions of the mass profile of the bar and disk. We find a general agreement between the bar parameters measured from our framework and literature measurements that are based on similar methods (e.g. Jiménez-Arranz et al., 2024b).
@@
Himansh Rathore would like to acknowledge interesting discussions with Mathieu Renzo, Peter Behroozi, Leandro Beraldo-e-Silva, David Sand, Dennis Zaritsky, Ó. Jiménez-Arranz, Ewa L. Łokas, Anthony Brown, Alfred Castro-Ginard, Tristan Cantat-Gaudin, Lia Athanassoula, Roeland van der Marel and Kathryne J. Daniel regarding the LMC’s bar as well as the completeness calibration framework. We would like to thank the anonymous referee for providing insightful comments that improved the quality as well as clarity of the paper. We would like to thank the journal data editor for suggestions which helped to add and improve software citations in the manuscript. This work utilized the Puma and ElGato High Performance Computing clusters at the University of Arizona. Himansh Rathore and Gurtina Besla are supported by NSF CAREER AST 1941096, NASA ATP 80NSSC24K1225 and NASA FINESST 23-ASTRO23-0004. This research uses services or data provided by the Astro Data Lab, which is part of the Community Science and Data Center (CSDC) Program of NSF NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the U.S. National Science Foundation. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. We respectfully acknowledge that the University of Arizona is on the land and territories of Indigenous peoples. Today, Arizona is home to 22 federally recognized tribes, with Tucson being home to the O’odham and the Yaqui. Committed to diversity and inclusion, the University strives to build sustainable relationships with sovereign Native Nations and Indigenous communities through education offerings, partnerships, and community service.
@@ Appendix A Gaia Data Selection
where $\varpi$ is the parallax in mas; $\mu_{\alpha\ast}$ and $\mu_{\delta}$ are the proper motions in Right Ascension (RA; $\alpha$, $\alpha^{\ast}=\alpha\cos\delta$) and declination (DEC; $\delta$) in mas yr-1 respectively; $\rho$ is the galactic radius from the LMC kinematic center $(\alpha_{0},\delta_{0})=(80.443^{\circ},-69.272^{\circ})$ as determined in C22. We use the following query in the Gaia query interface:
@@ Appendix B Coordinate Systems
The LMC in-plane cartesian coordinates can be obtained by the following transformation: